Question 6 1 pts Intercoastal Corporation recently paid a common stock dividend of $2.50. Analysts predict that the firm and its dividends will grow at a rate of 12% per year. Intercoastal has a beta of 1.35. The current risk-free rate of interest is 5.0% and the current return on the market is 13%. What is the intrinsic value of Intercoastal common stock? O $114.50 O $73.68 O $100 $140.34 $127.20
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

ke Microsoft Excel Home nert Page Layout Formulas Data Review View dd-Ins s Cut ta copy. E AutoSum ー E ゴWrap Text в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard Font Alignment Number Styles Cells Edting 296 289 290 291 292 293 294 295 296 297 298 299 300 301 302 303 304 305 306 307 4HKE CAPM UTILITY, SH ke- Rf+beta(Rm-Rf) ke-5% + 1.35(13%-5%) D1 D0(1+g) 2.50(1+0.12)- PO = D1/(ke-g) = 2.80/(0.158-0.12)- 15.80% 2.80 73.68 INTRINSIC VALUE PO:73.68 beta bond c future INDEX INTL CAP BUD ING PV, FV, ANNUITYDIR cleanYIELD bond stru WACC ex div DN11 erences: BD109 130% 16-10-2018